OpenStudy (anonymous):

The perfect ideal gas law... PV=nRT, can anyone give a question to this and how to apply this ideal perfect gas law to it??

OpenStudy (anonymous):

go to the link i gave you earlier and i can explain through there.

OpenStudy (anonymous):

Well so far what do you know about PV = nRT?

OpenStudy (anonymous):

aaaah.. sure. im in now

OpenStudy (anonymous):

hi k let say im okay withit but i need a harder example..

OpenStudy (anonymous):

oh... and i saw what ecah1 asked.. and yes im okay with that question but need some other tricky ones :D

OpenStudy (anonymous):

Harder example eh? I think "harder" meaning it could only be hard if you have to change units

OpenStudy (anonymous):

kind of.

OpenStudy (anonymous):

did you get it from the link? it on the lhs

OpenStudy (anonymous):

thank you... i see it now. that helps a lot. aligato

OpenStudy (anonymous):

sweet

OpenStudy (anonymous):

Or if you really want a difficult here you go: 0.56 g of Ethene gas is contained in a vessel at a pressure of 102KPa and Temperature is 30 C, Calculate the volume give the Answer is Cm^3

OpenStudy (anonymous):

Try doing this

OpenStudy (anonymous):

wat the... this is hard for me..lol.. fantasio is off atm.. he must be dizzy lookin at this :D in a good way :D

OpenStudy (anonymous):

oh and can you help me with the other question i asked?? if u have time tho

Parth (parthkohli):

\[\rm pressure \times volume = number ~of~moles \times temperature \times gas ~constant\]You only have to plug the information give to you in this. That's it. In almost all the problems, there would be only one unknown and you'll have to solve for that given the information.

Parth (parthkohli):

given to you*

OpenStudy (anonymous):

Agreed

OpenStudy (somy):

@ParthKohli units are important too and also n should be replaced by mass/Mr

Parth (parthkohli):

all of the units are accommodated in the gas constant, so yes, you'll have to convert the units to the ones that get cancelled out in \(R\). we can check all that stuff. just for instance, if you have \(\rm force = mass \times some~ constant\) and the constant is given to have the units \(m/s^2\), and force is given to have the units \(g~m/s^2\), then I'd write the mass in grams.

OpenStudy (anonymous):

yes I think the solver should have the both knowledge of how to cancel out units AND know where to plug in the numbers too so yea both of you are right :D

OpenStudy (anonymous):

thank you guys for all of the response